OpenStudy (anonymous):

Simplify the following expression: (x - 8y) - (2x - 4y). If the final answer is written in the form Ax + By, what is the value of B?

OpenStudy (anonymous):

x-8y-2x+4y = -x-4y A=-1,B=-4

OpenStudy (hoblos):

(x - 8y) - (2x - 4y) = x-8y -2x+4y = -x -4y B=-4

OpenStudy (anonymous):

Thanks people- can you help me on this question. Simplify the following expression: (5x^2 + 3x + 4) - (2x^2 + 5x - 1). If the final answer is written in the form Ax^2 + Bx + C, what is the value of A?

OpenStudy (anonymous):

^ its 9.